# GenUP: Generative User Profilers as In-Context Learners for Next POI Recommender Systems

### v2

@@ Abstract.
Traditional Point-of-Interest (POI) recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings. Furthermore, the cold-start problem—where systems have insufficient data for new users—limits their ability to generate accurate recommendations. Existing methods often address this by leveraging similar trajectories from other users, but this approach can be computationally expensive and increases the context length for LLM-based methods, making them difficult to scale. To address these limitations, we propose a method that generates natural language (NL) user profiles from large-scale, location-based social network (LBSN) check-ins, utilizing robust personality assessments and behavioral theories. These NL profiles capture user preferences, routines, and behaviors, improving POI prediction accuracy while offering enhanced transparency. By incorporating NL profiles as system prompts to LLMs, our approach reduces reliance on extensive historical data, while remaining flexible, easily updated, and computationally efficient. Our method is not only competitive with other LLM-based and complex agentic frameworks but is also more scalable for real-world POI recommender systems. Results demonstrate that our approach consistently outperforms baseline methods, offering a more interpretable and resource-efficient solution for POI recommendation systems. Our source code is available at: https://github.com/w11wo/GenUP/.
@@ 1. Introduction
Another challenge in POI recommendation is personalization, often achieved through dense vector representations of users. While these embeddings effectively model user behavior, they lack interpretability, transparency, and scrutability—qualities crucial for building user trust and improving recommendation explainability (O’Donovan and Smyth, 2005; Zhang and Chen, 2020). Natural language (NL) user profiles offer a promising alternative, where new users can quickly provide a concise description of their preferences, and existing users can update their profiles for more personalized and transparent recommendations (Radlinski et al., 2022). Despite their potential, NL profiles have not seen widespread adoption due to difficulties in evaluation and the ongoing development of methods for accurately generating these profiles (Ramos et al., 2024). Prior work, such as LLMob (Wang et al., 2024a), relies on rigid user categorization, generating multiple candidate personas and selecting the most likely one based on a given trajectory. While effective, this approach limits personalization by constraining users to predefined categories.
In response, we introduce NL user profiles for next POI recommendation, replacing long historical trajectory prompts with concise profiles to reduce computational overhead and context length. We propose generating these profiles from LBSN check-ins to learn user behaviors and integrate robust behavioral theories for more accurate POI predictions. Our contributions are as follows:
We introduce natural language (NL) user profiles as LLM system prompts for next POI recommendation.
We reconstruct NL user profiles from LBSN data, capturing individualized routines without predefined categories.
We combine LLM generation and behavioral theories for more grounded user profiles.
Our method enhances POI recommendation by improving transparency, reducing reliance on extensive historical data, and maintaining competitive accuracy with a more scalable, resource-efficient design. Unlike existing LLM-based approaches, it achieves strong performance while using smaller, open-source models, making it more adaptable for real-world deployment.
@@ 4.1. Natural Language User Profile Generation
Following the approach of (Li et al., 2024) and using their prompt formatting structure, user check-in records and trajectories were first transformed into prompts. Trajectory prompting (Radford et al., 2021; Xue et al., 2022; Zhou et al., 2022) allows the conversion of heterogeneous data from LBSNs into natural language sentences that LLMs can understand. These trajectory prompts were then concatenated and used to generate NL user profiles, as outlined in Table 1. In addition to generating a 200-word NL user profile, the LLM was tasked with predicting the 3 components above. To facilitate easier parsing, the output is structured in JSON format. Examples of a user profile can be found in Appendix A.
Validation and Challenges
A key challenge in user profile generation is validation, as extensively discussed in §2.2. Another challenge we encountered was LLM hallucinations, where it failed to follow the instructed JSON schema, leading to inconsistent parsing due to changed JSON key names. However, this issue affected fewer than five users and was easily resolved by re-running the prompt.
To validate our generated profiles, we ensured that predicted personality traits were valid BFI categories and that user attributes matched predefined classes. Additionally, we employed the LLM-as-a-Judge method (Zheng et al., 2023), where LLMs assessed the quality of generated profiles. While this approach has potential biases, we consider it a reasonable proxy for validation and have provided further details in Appendix B. Interestingly, our user profiles are also weak in conciseness, parallel with the findings of (Ramos et al., 2024), who employed human evaluators. To further evaluate the biases in our generated NL user profiles, we present their distribution in Appendix C.
@@ 4.2. Trajectory Prompting with Natural Language User Profiles
The current trajectory block consists of all check-in records from the current session, except for the latest record, which is placed in the target block. Unlike LLM4POI, we excluded historical check-in records from the user and other users to isolate the impact of our proposed NL user profiles and minimize the input context length.222Our method only implicitly incorporates historical information into its user profile, unlike LLM4POI. The latter has access to multiple similar, repeatable trajectories, whereas the former can only summarize frequent or long-term trajectories into a concise user profile. LLM4POI also incorporates intra-user social data, while ours does not. The range of POI IDs specified in the instruction block helps guide the LLM in generating a predicted POI ID, ensuring it falls within the set of POIs . The target block, which includes the next timestamp, user ID, and next POI ID, serves as the generation targets/labels and is excluded from the input during the prediction process.
@@ 5.1.1. Datasets
As we aimed to compare our approach with that of LLM4POI, we used the same public datasets they used and employed their pre-processing code333https://github.com/neolifer/LLM4POI. Specifically, we used the New York City (NYC) and Tokyo (TKY) Check-in Dataset (Yang et al., 2015b), which contains user check-ins from 12 April 2012 to 16 February 2013 from the popular LBSN, FourSquare444https://foursquare.com/. They also used the Gowalla-CA dataset (Yan et al., 2023), a subset of the original Gowalla555https://www.gowalla.com dataset (Cho et al., 2011), and filtered for check-in records from California and Nevada using geometric coordinates. Following their procedures, the filtered and pre-processed records are divided into training, validation, and test sets, using the first 80% for training, the middle 10% for validation, and the final 10% for testing. Only users and POIs that appear in the training set are kept in the validation and testing sets.
Furthermore, we recreated subsets of the Global-scale Check-in Dataset (Yang et al., 2019, 2022b), also derived from FourSquare, following the pre-processing methods outlined by (Feng et al., 2024). Trajectories were grouped into 72-hour sessions, with users having fewer than 5 sessions and sessions with fewer than 4 check-ins removed. The data was split into 70% for training, 10% for validation, and 20% for testing. Only users with between 3 and 50 sessions were included in the testing set. While they only used 200 test samples due to cost constraints, we evaluated on the entire testing set. We applied this pre-processing procedure to the cities of Moscow and São Paulo, two of the largest cities included in their evaluation. However, as shown in Appendix D, our reproduction of their dataset is not identical, as they filtered trajectories using an unspecified minimum distance from the city centers. In contrast, we used actual geometric boundaries of the cities obtained from OpenStreetMap666https://www.openstreetmap.org and included only the POIs located within those boundaries. All of the datasets we used are publicly available and anonymous.
@@ 5.1.2. Baselines
We compared our approach with the four kinds of baselines: supervised fine-tuned (SFT) LLMs, zero-shot in-context learning LLMs, agentic LLM frameworks, and graph-based models.
@@
AgentMove (Feng et al., 2024): Combines multiple LLMs as agents and integrates external knowledge sources such as urban structure, geospatial knowledge, and location graphs. Each agentic LLM serves as a different component, providing different information to the primary zero-shot, POI predictor LLM.
@@ 5.1.3. Evaluation
We evaluated our model using Accuracy@1, the same metric employed by both LLM4POI and AgentMove, to assess prediction accuracy when a user can only visit one next POI rather than multiple locations.
@@ 5.1.4. Models
To ensure a fair comparison against the reported results of LLM4POI (Li et al., 2024), we used the same pre-trained LLM, Llama-2-7b-longlora-32k (Chen et al., 2024a), for our main comparative experiments. In their setup, the model without historical trajectories is denoted as LLM4POI*, while the model trained with historical trajectories from both the current user and other users is called LLM4POI. We denote our models with the prefix GenUP (generative user profilers).
Moreover, we fine-tuned more recent variants of the LLaMA family of models (Touvron et al., 2023a; Dubey et al., 2024), including Llama 3.1 8B and Llama 3.2 1B, to further benchmark our method against newer architectures and evaluate their performance on the next POI prediction task. The inclusion of the 1 billion model variant allows for a more efficient inference while maintaining similar performances and requiring limited computational resources, making it more accessible for researchers with constrained hardware.
@@ 5.1.5. Implementation
Our fine-tuning implementation details are provided in Appendix E, based on the setup from (Li et al., 2024). We used a linearly decaying learning rate scheduler for all fine-tuning runs, with 20 warmup steps and no weight decay. Since we excluded historical trajectories from our prompts, we used a context length of 16,384 tokens and trained for 3 epochs. The Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020; Wolf et al., 2020) was used throughout our SFT experiments, which were conducted on NVIDIA H100 GPUs.
5.2. Results
@@
The evaluation results from the NYC, TKY, and CA datasets demonstrate the effectiveness of our proposed method, which consistently outperforms LLM4POI*, both of which utilize the same Llama 2 base LLM. Specifically, GenUP-Llama-2-7B achieves higher accuracy in NYC (0.2575 vs. 0.2356), TKY (0.1699 vs. 0.1517), and CA (0.1094 vs. 0.1016). These findings indicate that incorporating natural language user profiles as system prompts improves performance, even when using the same pre-trained LLM.
Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all three cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339). Notably, this model demonstrated competitive results compared to GETNext which exploits collaborative signals from cross-user data. Similarly, the smaller GenUP-Llama-3.2-1B model also consistently surpasses the LLM4POI* baseline, showing an improvement in accuracy despite using a smaller yet more powerful pre-trained base LLM.
@@
| LLM4POI* | Llama 2 7B | 0.146 | 0.166 |
| GenUP-Llama-2-7B | Llama 2 7B | 0.159 | 0.175 |
| GenUP-Llama-3.1-8B | Llama 3.1 8B | 0.163 | 0.178 |
| GenUP-Llama-3.2-1B | Llama 3.2 1B | 0.161 | 0.175 |
| In-context learning | |||
| LLM-Mob† | GPT-4o Mini | 0.080 | 0.140 |
| LLM-ZS† | GPT-4o Mini | 0.120 | 0.165 |
| Agentic LLM with external knowledge | |||
| AgentMove (base)† | GPT-4o Mini | 0.120 | 0.165 |
| AgentMove (+memory)† | GPT-4o Mini | 0.160 | 0.190 |
| AgentMove (+world)† | GPT-4o Mini | 0.090 | 0.175 |
| AgentMove (+collective)† | GPT-4o Mini | 0.140 | 0.175 |
| AgentMove (+mem+world)† | GPT-4o Mini | 0.135 | 0.240 |
| AgentMove (all)† | GPT-4o Mini | 0.150 | 0.230 |
We also compare the results of our proposed method with the LLM-based agentic framework, such as AgentMove, which utilizes multiple LLMs as agents alongside external knowledge sources. Additionally, we include their reported results for other LLM-based baselines that rely on historical trajectories or employ in-context learning rather than SFT. We present our results in Table 4.
Although our dataset reproduction is not identical to that of (Feng et al., 2024), our models achieved competitive results on Moscow and São Paulo datasets compared to those that leverage in-context learning (ICL) methods, such as LLM-Mob and LLM-ZS. Both ICL approaches incorporate historical trajectories to inform their predictions, while AgentMove combines multiple LLM agents with external knowledge sources to enhance its predictive performance. In contrast, our method solely employs a single LLM base with supervised fine-tuning (SFT), highlighting the effectiveness of our approach even with a more straightforward architecture.
Notably, our models remain competitive with AgentMove, despite the latter (1) utilizing a much larger GPT-4o Mini model, (2) employing multiple LLM agents, and (3) incorporating external knowledge sources, such as a world knowledge generator and a collective knowledge extractor from cross-user mobility data. This further underscores the robustness of our approach, which achieves comparable performance with significantly fewer resources.
Likewise, our GenUP-Llama-2-7B model continues to outperform the LLM4POI* baseline across both datasets, while the GenUP-Llama-3.1-8B remains the top performer overall, even surpassing AgentMove with its extensive external knowledge components on the Moscow dataset. Additionally, the GenUP-Llama-3.2-1B model remains competitive with its larger counterparts, demonstrating comparable performance despite its reduced size.

### v3

@@ Abstract.
Traditional Point-of-Interest (POI) recommendation systems often lack transparency, interpretability, and scrutability due to their reliance on dense vector-based user embeddings. Furthermore, the cold-start problem—where systems have insufficient data for new users—limits their ability to generate accurate recommendations. Existing methods often address this by leveraging similar trajectories from other users, but this approach can be computationally expensive and increases the context length for LLM-based methods, making them difficult to scale. To address these limitations, we propose a method that generates natural language (NL) user profiles from large-scale, location-based social network (LBSN) check-ins, utilizing robust personality assessments and behavioral theories. These NL profiles capture user preferences, routines, and behaviors, improving POI prediction accuracy while offering enhanced transparency. By incorporating NL profiles as system prompts to LLMs, our approach reduces reliance on extensive historical data, while remaining flexible, easily updated, and computationally efficient. Our method is not only competitive with other LLM-based methods but is also more scalable for real-world POI recommender systems. Results demonstrate that our approach consistently outperforms baseline methods, offering a more interpretable and resource-efficient solution for POI recommendation systems. Our source code is available at: https://github.com/w11wo/GenUP/.
@@ 1. Introduction
Another challenge in POI recommendation is personalization, often achieved through dense vector representations of users. While these embeddings effectively model user behavior, they lack interpretability, transparency, and scrutability—qualities crucial for building user trust and improving recommendation explainability (O’Donovan and Smyth, 2005; Zhang and Chen, 2020). Natural language (NL) user profiles offer a promising alternative, where new users can quickly provide a concise description of their preferences, and existing users can update their profiles for more personalized and transparent recommendations (Radlinski et al., 2022). Despite their potential, NL profiles have not seen widespread adoption due to difficulties in evaluation and the ongoing development of methods for accurately generating these profiles (Ramos et al., 2024). Prior work, such as LLMob (Wang et al., 2024a), relies on rigid user categorization, generating multiple candidate personas and selecting the most likely one based on a given trajectory. While effective, this approach limits personalization by constraining users to predefined categories. In response, we introduce NL user profiles for next POI recommendation, replacing long historical trajectory prompts with concise profiles to reduce computational overhead and context length. We propose generating these profiles from LBSN check-ins to learn user behaviors and integrate robust behavioral theories for more accurate POI predictions.
Contributions
We introduce a novel approach for deriving NL user profiles from implicit spatiotemporal signals in check-in data, a domain that has been largely overlooked in prior work. Existing LLM-driven recommender systems predominantly operate on datasets with explicit user feedback (e.g., ratings, reviews), whereas our method infers user preferences and routines from raw LBSN trajectories, requiring higher-level behavioral inference. To the best of our knowledge, we are the first to generate NL user profiles directly from LBSN check-in sequences and integrate them into a next POI recommendation framework. Our approach constructs these profiles using structured prompting that combines user attributes and personality traits, specifically leveraging components from the Theory of Planned Behavior and the Big Five Inventory. Unlike prior prompt-based user modeling methods, which apply generic persona templates or fixed category assignments, our profiles are data-driven, dynamically generated, and domain-specific, tailored to spatiotemporal mobility patterns. Our method enhances LLM-based POI recommendation by improving transparency, reducing reliance on extensive historical data, and maintaining competitive accuracy with a more scalable, resource-efficient design.
@@ 4.1. Natural Language User Profile Generation
Following the approach of (Li et al., 2024) and using their prompt formatting structure, user check-in records and trajectories were first transformed into prompts. Trajectory prompting (Radford et al., 2021; Xue et al., 2022; Zhou et al., 2022) allows the conversion of heterogeneous data from LBSNs into natural language sentences that LLMs can understand. These trajectory prompts were then concatenated and used to generate NL user profiles, as outlined in Table 1. In addition to generating a 200-word NL user profile, the LLM was tasked with predicting the 3 components above. To facilitate parsing, the output is returned in JSON format.
4.1.1. User Profile Validation and Challenges
A key challenge in user profile generation is validation, as extensively discussed in §2.2. Another challenge we encountered was LLM hallucinations, where it failed to follow the instructed JSON schema, leading to inconsistent parsing due to changed JSON key names. However, this issue affected fewer than five users and was easily resolved by re-running the prompt. To validate our generated profiles, we first ensured that predicted personality traits were valid BFI categories and that user attributes matched predefined classes.
Additionally, we employed the LLM-as-a-Judge method (Zheng et al., 2023), where LLMs assessed the quality of generated profiles. Following the qualitative evaluation approach proposed by (Ramos et al., 2024), we adapted their fluency, informativeness, conciseness, and relevance metrics to assess our NL profiles. LLMs were prompted using the metrics shown in Table 2, and their evaluation results are shown in Table 3. To mitigate bias from a single model’s judgment, we conducted evaluations using two LLMs: GPT-4o Mini and Gemini 1.5 Flash. Notably, conciseness emerged as a consistent weakness across both judges, mirroring observations from (Ramos et al., 2024), where human annotators similarly flagged verbosity in LLM-generated profiles. However, on the other metrics, our profiles were rated favorably, indicating that despite verbosity, they remain fluent, informative, and relevant.
To further investigate potential biases in our generated NL user profiles, we analyzed their distributions across three datasets: NYC, CA, and TKY. Table 4 presents the distribution of predicted BFI traits, while Table 5 summarizes the user attribute distributions. These statistics highlight the diversity of inferred user characteristics within and across datasets, providing insight into the diversity and variability of the generated profiles.
@@ 4.1.1. User Profile Validation and Challenges
| Given the following natural language (NL) user profile [user profile], you are to assess the user profile based on four criteria: |
| - Fluency: Is the NL profile both syntactically and semantically correct? |
| - Informativeness: Does the NL profile provide important information for a user profile? |
| - Conciseness: Is the NL profile written in a concise manner? |
| - Relevance: Given the list of reviews, is the NL profile relevant to the user? |
| Metric | NYC | TKY | CA |
| Judge: GPT 4-o Mini | |||
| Fluency | 100.00 | 100.00 | 100.00 |
| Informativeness | 100.00 | 100.00 | 100.00 |
| Conciseness | 43.17 | 40.60 | 37.26 |
| Relevance | 100.00 | 100.00 | 100.00 |
| Judge: Gemini 1.5 Flash | |||
| Fluency | 100.00 | 100.00 | 100.00 |
| Informativeness | 100.00 | 100.00 | 100.00 |
| Conciseness | 65.81 | 63.48 | 57.68 |
| Relevance | 100.00 | 100.00 | 100.00 |
@@ 4.2. Trajectory Prompting with Natural Language User Profiles
The current trajectory block consists of all check-in records from the current session, except for the latest record, which is placed in the target block. Unlike LLM4POI, we excluded historical check-in records from the user and other users to isolate the impact of our proposed NL user profiles and minimize the input context length.222Our method only implicitly incorporates historical information into its user profile, unlike LLM4POI. The latter has access to other similar trajectories, whereas the former can only summarize frequent or long-term trajectories into a concise user profile. LLM4POI also incorporates intra-user social data, while ours does not. The range of POI IDs specified in the instruction block helps guide the LLM in generating a predicted POI ID, ensuring it falls within the set of POIs . The target block, which includes the next timestamp, user ID, and next POI ID, serves as the generation targets/labels and is excluded from the input during the prediction process.
| Trait | NYC | TKY | CA |
| Extroverted | 736 | 1476 | 3526 |
| Introverted | 311 | 805 | 430 |
| Agreeable | 1047 | 2281 | 3956 |
| Antagonistic | 0 | 0 | 0 |
| Conscientious | 1047 | 2281 | 3956 |
| Unconscientious | 0 | 0 | 0 |
| Neurotic | 41 | 99 | 268 |
| Emotionally Stable | 1006 | 2182 | 3688 |
| Open to Experience | 1043 | 2279 | 3947 |
| Closed to Experience | 4 | 2 | 9 |
| Attribute | NYC | TKY | CA | |
| Age | Child | 1 | 2 | 14 |
| Adolescent | 72 | 274 | 330 | |
| Adult | 974 | 2005 | 3612 | |
| Older Adult | 0 | 0 | 0 | |
| Gender | Male | 774 | 1881 | 2911 |
| Female | 273 | 400 | 1045 | |
| Education | Some Schooling | 43 | 192 | 190 |
| High School | 55 | 165 | 176 | |
| College & Beyond | 949 | 1924 | 3590 | |
| Socioeconomic | Lower | 3 | 3 | 7 |
| Middle | 1028 | 2240 | 3651 | |
| Upper | 16 | 38 | 298 |
@@ 5.1. Setup
Maps of the recreated Moscow and São Paulo subsets from the Global-scale Check-in Dataset.
@@ 5.1.1. Datasets
As we aimed to compare our approach with that of LLM4POI, we used the same public datasets they used and employed their pre-processing code333https://github.com/neolifer/LLM4POI. Specifically, we used the New York City (NYC) and Tokyo (TKY) Check-in Dataset (Yang et al., 2015b), which contains user check-ins from 12 April 2012 to 16 February 2013 from the popular LBSN, Foursquare. They also used the Gowalla-CA dataset (Yan et al., 2023), a subset of the original Gowalla dataset (Cho et al., 2011), and filtered for check-in records from California and Nevada using geometric coordinates. Following their procedures, the filtered and pre-processed records are divided into training, validation, and test sets, using the first 80% for training, the middle 10% for validation, and the final 10% for testing. Only users and POIs that appear in the training set are kept in the validation and testing sets.
Furthermore, we recreated Moscow and São Paulo subsets of the Global-scale Check-in Dataset (Yang et al., 2019, 2022b), following the pre-processing methods outlined by (Feng et al., 2024). Unlike their approach, we used actual geometric boundaries of the cities obtained from OpenStreetMap444https://www.openstreetmap.org/ and included only the POIs located within those boundaries. Trajectories were grouped into 72-hour sessions, with users having fewer than 5 sessions and sessions with fewer than 4 check-ins removed. The data was split into 70% for training, 10% for validation, and 20% for testing. Only users with between 3 and 50 sessions were included in the testing set. Like (Feng et al., 2024), we limited the test set to 200 samples during evaluation due to cost constraints, but fine-tuned our models on the entire training set. All of the datasets we used are publicly available and anonymous. The maps of geometric boundaries of the two cities are shown in Figure 2(b).
@@ 5.1.2. Baselines
We compared our approach with the three kinds of baselines: supervised fine-tuned (SFT) LLMs, LLM-based in-context learning methods, and graph-based models.
@@ 5.1.3. Evaluation
We evaluated our model using Accuracy@1, aligning with the evaluation protocol of LLM4POI to ensure fair and consistent comparison. LLM4POI motivates the use of Accuracy@1 by framing next POI recommendation as a question-answering task, where only a single, contextually appropriate destination is returned, reflecting realistic constraints such as time-sensitive travel scenarios where only one top recommendation is allowed.
@@ 5.1.4. Models
To ensure a fair comparison against the reported results of LLM4POI (Li et al., 2024), we used the same pre-trained LLM, Llama-2-7b-longlora-32k (Chen et al., 2024a), for our main comparative experiments. In their setup, the model without historical trajectories is denoted as LLM4POI*, while the model trained with historical trajectories from both the current user and other users is called LLM4POI. We denote our models with the prefix GenUP (generative user profilers). Moreover, we fine-tuned more recent variants of the LLaMA family of models (Touvron et al., 2023a; Dubey et al., 2024), including Llama 3.1 8B and Llama 3.2 1B, to further benchmark our method against newer architectures and evaluate their performance on the next POI prediction task.
@@ 5.1.5. Implementation
Our fine-tuning implementation follows the setup described in (Li et al., 2024). We used a linearly decaying learning rate scheduler for all fine-tuning runs, with 20 warmup steps and no weight decay. Since we excluded historical trajectories from our prompts, we used a context length of 16,384 tokens and trained for 3 epochs. The Transformer Reinforcement Learning (TRL) library (von Werra et al., 2020; Wolf et al., 2020) was used throughout our SFT experiments, which were conducted on NVIDIA H100 GPUs.
5.2. Main Results
@@
The evaluation results from the NYC, TKY, and CA datasets demonstrate the effectiveness of our proposed method, which consistently outperforms LLM4POI*, both of which utilize the same Llama 2 base LLM. Specifically, GenUP-Llama-2-7B achieves higher accuracy in NYC (0.2575 vs. 0.2356), TKY (0.1699 vs. 0.1517), and CA (0.1094 vs. 0.1016). These findings indicate that incorporating natural language user profiles as system prompts improves performance, even when using the same pre-trained LLM. Moreover, the larger GenUP-Llama-3.1-8B model, further increases performance, achieving the highest accuracy in all three cities (NYC: 0.2582, TKY: 0.2127, CA: 0.1339). Notably, this model demonstrated competitive results compared to GETNext which exploits collaborative signals from cross-user data. Similarly, the smaller GenUP-Llama-3.2-1B model also consistently surpasses the LLM4POI* baseline, showing an improvement in accuracy despite using a smaller yet more powerful pre-trained base LLM.
We also compare the results of our proposed method with LLM-based baselines that rely on historical trajectories and employ in-context learning (ICL) rather than supervised fine-tuning (SFT). To ensure a fair comparison, we re-implemented LLM-Mob and LLM-ZS using Llama 3.1 8B Instruct to match the capabilities of base models. We present the evaluation results in Table 8. Our models consistently outperform LLM-Mob and LLM-ZS on both the Moscow and São Paulo datasets, despite those baselines leveraging ICL with historical trajectories as input. In contrast, GenUP uses compact NL user profiles to implicitly learn historical trajectories without relying on long context windows.
@@
| In-context learning with historical data | |||
| LLM-ZS | Llama 3.1 8B Inst. | 0.115 | 0.130 |
| LLM-Mob | Llama 3.1 8B Inst. | 0.145 | 0.120 |
@@
| GenUP-Llama-2-7B | Llama 2 7B | 0.180 | 0.205 |
| GenUP-Llama-3.1-8B | Llama 3.1 8B | 0.170 | 0.205 |
| GenUP-Llama-3.2-1B | Llama 3.2 1B | 0.180 | 0.205 |